You invest $7200 in two accounts paying 8% and 10% annual interest. At the end of the year the accounts earn the same interest. How much interest at each rate?
Answer by stanbon(75887)

You invest $7200 in two accounts paying 8% and 10% annual interest. At the end of the year the accounts earn the same interest. How much interest at each rate?
---
Equation:
0.08x = 0.10(7200-x)
Multiply both sides by 100 to get:
8x = 10(7200-x)
8x = 72000-10x
18x = 72000
x = 4000 (amount invested at 8%)
7200-x = 6800 (amount invested at 10%)
===========================================
Cheers,
Stan H.